When the price of a particular flavor of Pizza is tk. 400, buyer want to purchase 8000 units per year, but when the price falls to tk. 260 the quantity demanded rises to 10000 Pizza’s per year. Calculate the price elasticity of demand for Pizza and make comments your result. 


1
Expert's answer
2021-06-24T12:29:36-0400

"%change in quantity demanded =(8000-10000)*100\/8000=-25" %change in quantity demanded "%change in quantity demanded""= \frac{(8000-10000)×100}{8000}=-25"

% change in price "=\frac{(400-260)×100}{400}=35"

Price elasticity"= \frac{-25}{35}"

"Price elasticity = -0.7143"